OpenStudy (anonymous):

Solve 2^x - 8x + 10 = 0,x

OpenStudy (mathstudent55):

First, divide both sides by 2. Then try factoring or using the quadratic equation.

OpenStudy (shadowlegendx):

\[2^{x} - 8x + 10 = 0,x\]

OpenStudy (shadowlegendx):

What does 0,x mean?

OpenStudy (shadowlegendx):

The points on a graph?

OpenStudy (anonymous):

Solve for x

OpenStudy (shadowlegendx):

Yes I know it's solving for x but I don't understand the last part of that equation

OpenStudy (shadowlegendx):

@mathstudent55 can you shed some light on this?

OpenStudy (mathstudent55):

The equation ends in 0. The x means solve for x.

OpenStudy (shadowlegendx):

Oh, that ",x) confused me

OpenStudy (mathstudent55):

Think of the instructions as being: Solve each equation for the indicated variable: \(3x = 6, x\)

OpenStudy (shadowlegendx):

Yeah I usually don't see equations in that format

OpenStudy (mathstudent55):

I've come across a few like this, that 's how I figured it out.

OpenStudy (mathstudent55):

Sorry, I just noticed the exponent x. Forget what I wrote above. I misread it.

OpenStudy (shadowlegendx):

How would we divide by 0 though? We would get undefined

OpenStudy (shadowlegendx):

divide 0 by 2

OpenStudy (shadowlegendx):

Yeah c;

OpenStudy (shadowlegendx):

\[2^{x} - 10 = 8x\] Add 8x to both sides = 1st step?

OpenStudy (shadowlegendx):

@mathstudent55

OpenStudy (anonymous):

Yeah and its plus 10!

OpenStudy (anonymous):

You can use rules of logarithms to get rid of x from the exponent.

OpenStudy (mathstudent55):

First, 0/2 is not undefined. 0/2 = 0 2/0 is undefined. Second, since this equation has an exponent of x, forget all my instructions above, including dividing both sides by 2. I misread the first term as 2x^2.

OpenStudy (anonymous):

\[\ln(C^x) = xln(C)\]

OpenStudy (mathstudent55):

The problem with taking logs is that you will end you with a sum or difference involving a variable, and we can't simplify such a log. I only know how to solve this by trying to graph it and then by trial and error to see where the zeros are.

OpenStudy (anonymous):

Forgot about that 8x.

OpenStudy (anonymous):

Are you sure it is not\[x^2\] @Swqaedf

OpenStudy (anonymous):

Yes. It's 2^x

OpenStudy (campbell_st):

one solution is to graph the curve and identify the zeros... or use an approximation method... using a table of values the zeros and between 1 and 2, then 4 and 5 hope it helps